# Venkata Ramana Makkapati Honda Aircraft Company

### v1

@@ 4.1 Theoretical Results
$\implies t=\dfrac{a}{\sqrt{v_{D}^{2}-v_{I}^{2}}}+\dfrac{t_{s}}{\sqrt{v_{D}^{2}-v_{I}^{2}}}\Bigg[v_{I}\cos\theta$
$~~~~~~~~~~~~~~+\sqrt{v_{D}^{2}-v_{I}^{2}}-\sqrt{v_{D}^{2}-v_{I}^{2}\sin^{2}\theta}\Bigg]$ (20)
Since $t_{s}\geq 0$, $v_{D}>v_{I}$ and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, (20) yields the capture time $t\geq a/(\sqrt{v_{D}^{2}-v_{I}^{2}})=t_{2}$. ∎
As discussed in [makkapati2019optimal], the constant bearing strategy and the Apollonius circle have close connections. The Apollonius circle is defined for a pursuer-evader pair as the set of capture points when the pursuer follows the constant bearing strategy, and the evader chooses a constant heading angle. The geometric properties of the Apollonius circle were discussed in [ramana2017pursuit]. The center and radius of the Apolloinus circle can be obtained as $\left(av_{D}^{2}/(v^{2}_{D}-v_{I}^{2}),0\right)$ and $av_{D}v_{I}/v_{D}^{2}-v_{I}^{2}$, respectively.
It has been proven that the line segment joining the two tangent points $A_{1}$ and $A_{2}$, in Fig. 2, passes through the initial position of the independent agent [ramana2017pursuit]. Furthermore, the line joining $A_{1}$ and $A_{2}$ subtends an angle of $2\sin^{-1}(v_{I}/v_{D})$ at the initial position of the dependent agent, and the line is perpendicular to the initial LOS. At the same time, it is also evident that the line joining the points $P_{1}$ and $P_{2}$ subtends an angle of $2\sin^{-1}(v_{I}/v_{D})$ at the initial position of the dependent agent, and the line is perpendicular to the LOS. Therefore, the points $A_{1}$, $P_{1}$, and the initial position of the dependent agent are collinear. Similarly, the points $A_{2}$, $P_{2}$, and the initial position of the dependent agent are collinear. It can be noted that at time $t_{2}$, the initial position of the independent agent $x^{0}_{I}$ lies on the line segment $\overline{P_{1}P_{2}}$. Consequently, $\overline{P_{1}P_{2}}$ coincides with line segment $\overline{A_{1}A_{2}}$ at time $t_{2}$.
@@
The above corollary provides a bound for the DRS in the second scenario ($t_{2}<t\leq t_{c}$). The bound is characterized by the points $P_{1}$ and $P_{2}$, and is depicted in Fig. 4 using a bold line. It is important to note that the above corollary, while providing a tight bound for the DRS, does not define the boundary of the DRS. The mathematical proof to establish the boundary of the DRS for $t_{2}<t\leq t_{c}$ is elusive. While a formal mathematical proof remains an open challenge, we present compelling evidence using simulation results (in the following subsection) that supports our hypothesis below.
Hypothesis 4.6.
@@
$x\geq\dfrac{a^{2}+v_{D}^{2}t^{2}-v_{I}^{2}t^{2}}{2a}\Bigg\}.$ (21)
The vertical line $x=(a^{2}+v_{D}^{2}t^{2}-v_{I}^{2}t^{2})/2a$ contains the points $Q_{1}$ and $Q_{2}$, which are the intersection points for the boundaries of the reachable sets of both agents ($\partial\mathcal{R}_{I}$ and $\partial\mathcal{R}_{D}$), as shown in Fig. 4. In the following subsection, it is empirically shown that the DRS for time $t_{2}<t\leq t_{c}$ is the shaded region in Fig. 4, which is characterized by the minor segment $Q_{1}P_{x}Q_{2}$.
@@
The Limiting Case of $v_{D}=v_{I}$
In the case of $v_{D}=v_{I}$, at time $t$, the points $P_{1}$ and $P_{2}$ are obtained as $(0,v_{I}t)$ and $(0,-v_{I}t)$, respectively. Therefore, the points $P_{1}$, $P_{2}$, and the initial point of the dependent agent are collinear. The line joining the three points forms the vertical diameter of the circle $\partial\mathcal{R}_{D}(t)$ (boundary of the dependent agent’s reachable set), which can be visualized in Fig. 5
Note that as $v_{I}\rightarrow v_{D}$, $t_{2}\rightarrow\infty$. Consequently, only the first scenario discussed in Section 4.1 is relevant for the limiting case, and the second scenario is irrelevant. Furthermore, it can be observed that the conclusion drawn in the proof of Theorem 4.4 is valid for the case of $v_{D}=v_{I}$. Consequently, the DRS is given by
@@
Figure 5: DRS for the limiting case $v_{D}=v_{I}$
@@
4.2 Simulation Results
This subsection presents simulation results based on discrete-time point-cloud propagation to observe the evolution of the dependent reachable sets for the time $0\leq t\leq t_{c}$. The initial positions of both agents are $x_{D}^{0}(0,0)$ and $x_{I}^{0}(1,0)$ and their speeds are set to $v_{D}=1$ and $v_{I}=0.5$ respectively. Simulations are carried out with the time-step $\Delta t=0.2$ until all the independent agents are captured. To generate a point cloud, we consider that at a given time instant, an independent agent chooses eighteen equally-spaced heading angles from $(-\pi,\pi]$. Consequently, at the $n^{th}$ time step, $18^{n}$ points are generated for the independent agent. At any time instant, every independent agent point has an associated dependent agent point obtained via the constant bearing strategy. For the simulation parameters mentioned above, $t_{1}=1.0$, $t_{2}\approx 1.155$ and $t_{c}=2.0s$.
Figure 6 presents the results obtained from the point cloud-based simulations. The blue color corresponds to the dependent agent, and the red color corresponds to the independent agent. The dashed circles indicate the boundaries of the agents’ reachable sets. The green line indicates the boundary of the DRS. The inverted triangles represent the initial positions of the two agents. The points in the red point cloud represent uncaptured independent agents, and the points in the blue point cloud represent the active dependent agents that are yet to capture their corresponding independent agents (refer to Section 2). As the simulation progresses, when a dependent agent captures an independent agent, that particular pair is eliminated from their respective point clouds. Figures 6(a) and 6(b) represent two instances of the scenario $0<t\leq t_{1}$, and Fig. 6(c) represents the case where $t\approx t_{2}$. Figures 6(d)-6(f) depict three instances of $t_{2}<t\leq t_{c}$.
@@ 4.2 Simulation Results
In Figs. 6(d) and 6(e), first, it can be observed that the minor segment $P_{1}P_{x}P_{2}$ contains points that are not part of the active region of the independent agent’s reachable set. Note that in the case of constant bearing, the relative distance between the agents reduces along the LOS. Therefore, the points in the minor segment $P_{1}P_{x}P_{2}$ that are not part of the independent agent’s reachable set cannot be part of the DRS. Furthermore, there are no blue points in the region between the line segments $\overline{P_{1}P_{2}}$ and $\overline{Q_{1}Q_{2}}$. It can be observed that the chord $\overline{Q_{1}Q_{2}}$ is part of the boundary of the blue point cloud when $t_{2}<t\leq t_{c}$, and the minor segment $Q_{1}P_{x}Q_{2}$ contains all the blue points when $t_{2}<t\leq t_{c}$. A more enhanced image of the interaction between the reachable sets is shown in the upper left corner of each figure. Simulations have been conducted for different parameters ($v_{D}$, $v_{I}$, etc.), which have not been included for brevity, and the results provided the same observations. Thus, it could be empirically stated that the DRS, initially represented by the segment $P_{1}P_{x}P_{2}$, reaches a maximum area at $t=t_{2}$ and subsequently shrinks to the segment $Q_{1}P_{x}Q_{2}$, which is contained within the segment $P_{1}P_{x}P_{2}$. Figure 6(f) ($t=t_{c}$) presents the case where the dependent agent’s reachable set fully encompasses the independent agent’s reachable set, and all the points from both red and blue point clouds ended in capture.
@@ 5 The Optimization Problem
This section formulates an optimization problem and provides empirical results pertaining to the constant bearing strategy, and the corresponding DRS. An approach to identifying the DRS is to analytically define the set of all the points that the dependent agent can reach using the constant bearing strategy for a given point from the independent agent’s reachable set. In this regard, consider a point $\mathbf{x}_{I}(t)=\mathbf{x}=[x,y]^{\top}\in\mathcal{R}_{I}(t)$, which the independent agent reaches at time $0\leq t\leq t_{c}$. Note that if $\|\mathbf{x}\|_{2}<v_{I}t$, there are infinitely many trajectories that the independent agent can take to reach the point $\mathbf{x}$. The goal is to find all the points that the dependent agent, which follows the constant bearing strategy, can reach for the given $\mathbf{x}$.
As formulated in Section 2, since the dependent agent follows the constant bearing strategy and given the initial conditions of both agents (origin and $(a,0)$), the vertical coordinates of both agents are the same for all time $t\geq 0$. Therefore, the vertical coordinate of the dependent agent at time $t$ is the same as the independent agent’s vertical coordinate, which is $y$. Consequently, we intend to find the minimum and the maximum horizontal coordinates that the dependent agent can reach at the time instant $t$. Since $\dot{y}_{D}(\tau)=\dot{y}_{I}(\tau)$ in this scenario, $\dot{x}_{D}(\tau)=\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}$. In this regard, consider the problem of obtaining the extrema for the functional
@@
$\dfrac{\partial L}{\partial x_{I}}-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(\dfrac{\partial L}{\partial\dot{x}_{I}}\right)=-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(2\dot{x}_{I}(\tau)\lambda(\tau)\right)=0,$ (27)
$\dfrac{\partial L}{\partial y_{I}}-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(\dfrac{\partial L}{\partial\dot{y}_{I}}\right)=-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\Bigg(-\dfrac{\dot{y}_{I}(\tau)}{\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}}$
$~~~~~~~~~~~~~~~~~~~~~~~~~+2\lambda(\tau)\dot{y}_{I}(\tau)\Bigg)=0.$ (28)
@@
Equations (27) and (28) cannot be solved further to obtain analytical expressions of the extrema. Therefore, simulation-based empirical studies were conducted to understand the nature of the extrema in the case where the independent agent follows a single-switching control function to reach the point $\mathbf{x}$. Note that in such cases, the locus of points at which the independent agent switches between constant control inputs is the ellipse
@@
The above hypotheses are demonstrated using simulation results. Figure 7 depicts four cases from the numerous simulations that were performed to analyze the extrema of the optimization problem in (23)-(25). The simulations were conducted with $\mathbf{x}_{I}(0)=[0,0]^{\top}$ (red square), $v_{D}=1$, $v_{I}=0.5$, $t=8$. The four points (blue, black, purple and green coloured inverted triangles) chosen within $\mathcal{R}_{I}(t)$ (interior of the red circle) represent a subset of the independent agent’s possible locations at time $t$. In each case, the corresponding coloured ellipse represents the locus of switch points $\mathcal{E}$, as given in (31), that the independent agent could execute with a single-switching control function to reach the selected point from $x_{I}(0)$. For each ellipse, the integral in (23) is evaluated for 360 equally spaced points (in terms of angular displacement) on the ellipse to identify the maximum and minimum points. On each ellipse in Fig. 7, the maxima points are indicated using filled circles, and the minima points are represented using empty circles. Hence, it can be observed that the maxima points always correspond to the points with the maximum and the minimum horizontal coordinates on the ellipse. Similarly, the minima points coincide with the points having the maximum and the minimum vertical coordinates on the ellipse. As $\mathbf{x}_{I}(t)$ varies between the origin and the boundary of $\mathcal{R}_{I}(t)$, the locus of the switch points changes from circular to elliptical with increasing eccentricity.
Figure 7: Simulation results that empirically substantiate Hypotheses 5.1 and 5.2.

### v2

@@ 4.1 Theoretical Results
$\implies$ $t=\dfrac{a}{\sqrt{v_{D}^{2}-v_{I}^{2}}}+\dfrac{t_{s}}{\sqrt{v_{D}^{2}-v_{I}^{2}}}\Bigg[v_{I}\cos\theta+\sqrt{v_{D}^{2}-v_{I}^{2}}-\sqrt{v_{D}^{2}-v_{I}^{2}\sin^{2}\theta}\Bigg]$
$~~=\dfrac{a}{\sqrt{v_{D}^{2}-v_{I}^{2}}}+\dfrac{t_{s}}{\sqrt{v_{D}^{2}-v_{I}^{2}}}\Bigg[v_{I}\cos\theta+\sqrt{v_{D}^{2}-v_{I}^{2}}-\sqrt{v_{D}^{2}-v_{I}^{2}+v_{I}^{2}\cos^{2}\theta}\Bigg].$ (21)
Since $t_{s}\geq 0$, $v_{D}>v_{I}$ and $\theta\in\left[-\dfrac{\pi}{2},\dfrac{\pi}{2}\right]$, (21) yields the capture time $t\geq a/\left(\sqrt{v_{D}^{2}-v_{I}^{2}}\right)=t_{2}$. Consequently, the dependent agent cannot capture the independent agent, which executes the control function in (18), using the constant bearing pursuit strategy before time $t_{2}$. Therefore, the dependent agent can reach all the points from the set in (17) via active pursuit trajectories. Hence proved. ∎
As discussed in [16], the constant bearing pursuit strategy and the Apollonius circle have close connections. The Apollonius circle is defined for a pursuer-evader pair as the set of capture points when the pursuer follows the constant bearing pursuit strategy, and the evader chooses a constant heading angle. The geometric properties of the Apollonius circle can be found in a prior study [19]. Given the initial positions of the dependent and independent agents as $(0,0)$ and $(a,0)$, the center and radius of the corresponding Apollonius circle $\mathcal{A}$ can be obtained as $\left(av_{D}^{2}/(v^{2}_{D}-v_{I}^{2}),0\right)$ and $av_{D}v_{I}/v_{D}^{2}-v_{I}^{2}$, respectively, as shown in Figure 2.
Lemma 4.5.
For $0\leq t\leq t_{2}$, the points $P_{1}$ and $P_{2}$ lie on either tangent lines of the Apollonius circle $\mathcal{A}$ emanating from the initial position of the dependent agent.
Proof.
The line segment joining the two tangent points $A_{1}$ and $A_{2}$, in Figure 2, was shown to pass through the initial position of the independent agent (see Lemma 3.2 in Ref. [19]). Furthermore, the line joining $A_{1}$ and $A_{2}$ subtends an angle of $2\sin^{-1}(v_{I}/v_{D})$ at the initial position of the dependent agent, and the line is perpendicular to the initial LOS (see Lemma 3.1 in Ref. [19]). Given the coordinates of $P_{1}$ and $P_{2}$ as $\left(t\sqrt{v_{D}^{2}-v_{I}^{2}},\pm v_{I}t\right)$, the line joining the origin and $P_{1}$ (or $P_{2}$) is angled at $\sin^{-1}(v_{I}/v_{D})$ from the horizontal axis. Consequently, the line joining the points $P_{1}$ and $P_{2}$ subtends an angle of $2\sin^{-1}(v_{I}/v_{D})$ at the initial position of the dependent agent, and the line is perpendicular to the LOS. Therefore, the points $A_{1}$, $P_{1}$, and the initial position of the dependent agent are collinear. Similarly, the points $A_{2}$, $P_{2}$, and the initial position of the dependent agent are collinear. Hence proved. ∎
It can be noted that at time $t_{2}$, the initial position of the independent agent $\mathbf{x}^{0}_{I}=(a,0)$ lies on the line segment $\overline{P_{1}P_{2}}$ [19]. Therefore, at time $t_{2}$, $\overline{P_{1}P_{2}}$ coincides with the line segment $\overline{A_{1}A_{2}}$ at time $t_{2}$.
@@
The above corollary provides a bound for the DRS in the second scenario ($t_{2}<t\leq t_{c}$). The bound is characterized by the points $P_{1}$ and $P_{2}$, and is depicted in Figure 4. It is important to note that the above corollary, while providing a tight bound for the DRS, does not define the boundary of the DRS. The mathematical proof to establish the boundary of the DRS for $t_{2}<t\leq t_{c}$ is elusive. While a formal mathematical proof remains an open challenge, we present compelling evidence using simulation results (in the following subsection) that supports our hypothesis below.
Hypothesis 4.7.
@@
The vertical line $x=(a^{2}+v_{D}^{2}t^{2}-v_{I}^{2}t^{2})/2a$ contains the points $Q_{1}$ and $Q_{2}$, which are the intersection points for the boundaries of the reachable sets of both agents ($\partial\mathcal{R}_{I}$ and $\partial\mathcal{R}_{D}$), as shown in Figure 4. In Section 4.3, it is empirically shown that the DRS for time $t_{2}<t\leq t_{c}$ is the shaded region in Figure 4, which is characterized by the minor segment $Q_{1}P_{x}Q_{2}$.
@@
4.2 The Limiting Case of $v_{D}=v_{I}$
In the case of $v_{D}=v_{I}$, at time $t$, the points $P_{1}$ and $P_{2}$ are obtained as $(0,v_{I}t)$ and $(0,-v_{I}t)$, respectively. Therefore, the points $P_{1}$, $P_{2}$, and the initial point of the dependent agent are collinear. The line joining the three points forms the vertical diameter of the circle $\partial\mathcal{R}_{D}(t)$ (boundary of the dependent agent’s reachable set), which can be visualized in Figure 5
Figure 5: DRS for the limiting case $v_{D}=v_{I}$
Note that as $v_{D}\rightarrow v_{I}$, $t_{2}\rightarrow\infty$. Consequently, only the first scenario discussed in Section 4.1 is relevant for the limiting case, and the second scenario is irrelevant. Furthermore, it can be observed that the conclusion drawn in the proof of Theorem 4.4 is valid for the case of $v_{D}=v_{I}$. Consequently, the DRS is given by
@@ 4.2 The Limiting Case of $v_{D}=v_{I}$
4.3 Simulation Results
This subsection presents simulation results based on discrete-time point-cloud propagation to observe the evolution of the dependent reachable sets over time $0\leq t\leq t_{c}$. While HJ reachability-based numerical methods provide computational efficiency in estimating reachable sets, the provision to track active trajectories is unavailable. Point-cloud propagation is chosen to accurately estimate the DRS by visualizing the active trajectories and removing those trajectories that result in capture. The initial positions of both agents are $x_{D}^{0}(0,0)$ and $x_{I}^{0}(1,0)$ and their speeds are set to $v_{D}=1$ and $v_{I}=0.5$ respectively. Simulations are carried out with the time-step $\Delta t=0.2$ until all the independent agents are captured. To generate a point cloud, we consider that at a given time instant, an independent agent chooses eighteen equally-spaced heading angles from $(-\pi,\pi]$. Consequently, at the $n^{th}$ time step, $18^{n}$ points are generated for the independent agent. At any time instant, every independent agent point has an associated dependent agent point obtained via the constant bearing pursuit strategy. For the simulation parameters mentioned above, $t_{1}=1.0$, $t_{2}\approx 1.155$ and $t_{c}=2.0s$.
@@ 4.3 Simulation Results
Figure 6 presents the results obtained from the point cloud-based simulations. The blue color corresponds to the dependent agent, and the red color corresponds to the independent agent. The dashed circles indicate the boundaries of the agents’ reachable sets. The green line indicates the boundary of the DRS. The inverted triangles represent the initial positions of the two agents. The points in the red point cloud represent uncaptured independent agents, and the points in the blue point cloud represent the active dependent agents that are yet to capture their corresponding independent agents (refer to Section 2). As the simulation progresses, when a dependent agent captures an independent agent, that particular pair is eliminated from their respective point clouds. Figures 6(a) and 6(b) represent two instances of the scenario $0<t\leq t_{1}$, and Figure 6(c) represents the case where $t\approx t_{2}$. Figures 6(d)-6(f) depict three instances of $t_{2}<t\leq t_{c}$.
@@
In Figs. 6(d) and 6(e), it can be observed that the minor segment $P_{1}P_{x}P_{2}$ contains a region that is outside of the independent agent’s reachable set $\mathcal{R}_{I}(t)$. Therefore, the points in the minor segment $P_{1}P_{x}P_{2}$ that are not part of the independent agent’s reachable set cannot be part of the DRS. Furthermore, there are no blue points in the region between the line segments $\overline{P_{1}P_{2}}$ and $\overline{Q_{1}Q_{2}}$. It can be observed that the chord $\overline{Q_{1}Q_{2}}$ is part of the boundary of the blue point cloud when $t_{2}<t\leq t_{c}$, and the minor segment $Q_{1}P_{x}Q_{2}$ contains all the blue points when $t_{2}<t\leq t_{c}$. A more enhanced image of the interaction between the reachable sets is shown in the upper left corner of each figure. Simulations have been conducted for different parameters ($v_{D}$, $v_{I}$, etc.), which have not been included for brevity, and the results provided the same observations. Thus, it could be empirically stated that the DRS, initially represented by the segment $P_{1}P_{x}P_{2}$, reaches a maximum area at $t=t_{2}$ and subsequently shrinks to the segment $Q_{1}P_{x}Q_{2}$, which is contained within the segment $P_{1}P_{x}P_{2}$. Figure 6(f) ($t=t_{c}$) presents the case where the dependent agent’s reachable set fully encompasses the independent agent’s reachable set, and all the points from both red and blue point clouds ended in capture.
@@ 5 The Optimization Problem
In this section, an optimization problem is formulated to gain deeper insight into the geometry of the DRS. A simple approach to determining the DRS at a time $t\geq 0$, $\mathcal{D}(t)$, consists of two steps. First, for each possible state $\mathbf{x}$ that the independent agent can reach at time $t$, compute the corresponding set of states $\mathcal{S}(\mathbf{x},t)$ that the dependent agent can reach with the constant bearing pursuit strategy. The complete DRS is then obtained as the union of all such sets of states of the dependent agent, i.e.
$\mathcal{D}(t)=\underset{\mathbf{x}\in\mathcal{R}_{I}(t)}{\bigcup}\mathcal{S}(\mathbf{x},t).$ (24)
As stated earlier in Section 1, it is important to note that the independent agent may reach a particular state at time $t$ via multiple distinct trajectories. In such cases, the dependent agent can potentially reach multiple different states from the same initial condition with the constant bearing pursuit strategy.
Now, consider a point $\mathbf{x}_{I}(t)=\mathbf{x}=[x,y]^{\top}\in\mathcal{R}_{I}(t)$, where the independent agent reaches at time $0\leq t\leq t_{c}$. Note that if $\|\mathbf{x}\|_{2}<v_{I}t$, there are infinitely many trajectories that the independent agent can take to reach the point $\mathbf{x}$. The goal is to find all the points that the dependent agent, which follows the constant bearing pursuit strategy, can reach for the given $\mathbf{x}$.
As formulated in Section 2, since the dependent agent follows the constant bearing pursuit strategy and given the initial conditions of both agents (origin and $(a,0)$), the vertical coordinates of both agents are the same for all time $t\geq 0$. Therefore, the vertical coordinate of the dependent agent at time $t$ is the same as the independent agent’s vertical coordinate, which is $y$. Subsequently, we intend to find the minimum and the maximum horizontal coordinates that the dependent agent can reach at the time instant $t$. Since $\dot{y}_{D}(\tau)=\dot{y}_{I}(\tau)$ for the constant bearing pursuit strategy, $\dot{x}_{D}(\tau)=\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}$. In this regard, consider the problem of obtaining the extrema for the functional
@@
$\dfrac{\partial L}{\partial x_{I}}-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(\dfrac{\partial L}{\partial\dot{x}_{I}}\right)=-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\Bigg(2\dot{x}_{I}(\tau)\lambda(\tau)\Bigg)=0,$ (29)
$\dfrac{\partial L}{\partial y_{I}}-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\left(\dfrac{\partial L}{\partial\dot{y}_{I}}\right)=-\dfrac{\mathrm{d}}{\mathrm{d}\tau}\Bigg(-\dfrac{\dot{y}_{I}(\tau)}{\sqrt{v_{D}^{2}-\dot{y}_{I}^{2}(\tau)}}+2\lambda(\tau)\dot{y}_{I}(\tau)\Bigg)=0.$ (30)
@@
Equations (29) and (30) cannot be solved further to obtain analytical expressions of the extrema. Therefore, simulation-based empirical studies were conducted to understand the nature of the extrema in the case where the independent agent follows a single-switching control function, which is the simplest candidate function, to reach the point $\mathbf{x}$. Note that in such cases, the locus of points at which the independent agent switches between constant control inputs is the ellipse
@@
The above hypotheses can be visualized using the simulation results shown in Figure 7, which depicts four distinct cases from the numerous simulations that were performed to analyze the extrema of the optimization problem in (25)-(27).
Figure 7: Simulation results that visualize the extrema of the optimization problem in (25)-(27), per Hypotheses 5.1 and 5.2.
The simulations corresponding to Figure 7 were conducted with $\mathbf{x}_{I}(0)=[0,0]^{\top}$ (red square), $v_{D}=1$, $v_{I}=0.5$, $t=8$. The four points (blue, black, purple, and green inverted triangles) chosen within $\mathcal{R}_{I}(t)$ (interior of the red circle) represent a subset of the possible locations of independent agent at time $t$. In each case, the corresponding colored ellipse represents the locus of switch points $\mathcal{E}$, as given in (33), that the independent agent could execute with a single-switching control function to reach the selected point from $x_{I}(0)$. For each ellipse, the integral in (25) is evaluated for 360 equally spaced points (in terms of angular displacement) on the ellipse to identify the maximum and minimum points. On each ellipse in Figure 7, the maxima points are indicated using filled circles, and the minima points are represented using empty circles. Hence, it can be observed that the maxima points always correspond to the points with the maximum and the minimum horizontal coordinates on the ellipse. Similarly, the minima points coincide with the points having the maximum and minimum vertical coordinates on the ellipse. As $\mathbf{x}_{I}(t)$ varies between the origin and the boundary of $\mathcal{R}_{I}(t)$, the locus of the switch points changes from circular to elliptical with increasing eccentricity.
Figure 8: Scatter plot representing the points from $\mathcal{R}_{I}(t)$ that are considered for Monte Carlo simulations.
To further validate Hypothesis 5.2, Monte Carlo simulations were conducted. To this end, the simulation parameters for the independent agent are fixed as $\mathbf{x}_{I}(0)=[0,0]^{\top}$, $v_{I}=1.2$, $t=5$. A total of 2,500 points are randomly chosen from the reachable set of the independent agent $\mathcal{R}_{I}(t)$, which is a circle with radius $6$, using the standard uniform distribution. A scatter plot of the random points along with $\partial\mathcal{R}_{I}(t)$ is shown in Figure 8. Four different speed ratios $v_{I}/v_{D}=0.8,0.6,0.4,0.2$, which correspond to $v_{D}=1.5,2,3,6$, respectively, are considered for the Monte Carlo simulations. Consequently, a total of 10,000 distinct ellipses are analyzed to identify the extrema of the integral in (25). For each ellipse, the extrema analytically computed using Hypothesis 5.2 is compared against the extrema that are obtained numerically. The maximum root mean square error for the four extremum points across the 10,000 ellipses is found to be of the order of $10^{-15}$, thus validating the hypothesis.